# Low dimensional representation of multi-patient flow cytometry datasets using optimal transport for minimal residual disease detection in leukemia

### v1

@@ Abstract
Representing and quantifying Minimal Residual Disease (MRD) in Acute Myeloid Leukemia (AML), a type of cancer that affects the blood and bone marrow, is essential in the prognosis and follow-up of AML patients. As traditional cytological analysis cannot detect leukemia cells below 5%, the analysis of flow cytometry dataset is expected to provide more reliable results. In this paper, we explore statistical learning methods based on optimal transport (OT) to achieve a relevant low-dimensional representation of multi-patient flow cytometry measurements (FCM) datasets considered as high-dimensional probability distributions. Using the framework of OT, we justify the use of the K-means algorithm for dimensionality reduction of multiple large-scale point clouds through mean measure quantization by merging all the data into a single point cloud. After this quantization step, the visualization of the intra and inter-patients FCM variability is carried out by embedding low-dimensional quantized probability measures into a linear space using either Wasserstein Principal Component Analysis (PCA) through linearized OT or log-ratio PCA of compositional data. Using a publicly available FCM dataset and a FCM dataset from Bordeaux University Hospital, we demonstrate the benefits of our approach over the popular kernel mean embedding technique for statistical learning from multiple high-dimensional probability distributions. We also highlight the usefulness of our methodology for low-dimensional projection and clustering patient measurements according to their level of MRD in AML from FCM. In particular, our OT-based approach allows a relevant and informative two-dimensional representation of the results of the FlowSom algorithm, a state-of-the-art method for the detection of MRD in AML using multi-patient FCM.
@@ 1 Introduction
Acute Myeloid Leukemia (AML) is a cancer of the blood and bone marrow when the development of blood stem cells (immature cells) into mature cells is altered, and the number of these immature cells, called myeloid blasts, increases. The diagnosis of AML consists in the detection of blasts which invade the body to great amounts, and it is typically assessed with the cytological analysis of the patient’s bone marrow. The goal of the chemotherapy treatment is therefore to eradicate the blasts population: 80% of patients reach complete remission after one cycle of chemotherapy, however 50% relapse [38]. The definition of complete remission states that the bone marrow should contain less than 5% of blasts. However, detecting smaller numbers of persisting residual leukemic cells, called Minimum Residual Disease (MRD), is critical for disease monitoring and relapse prognosis. This motivates the need to catch the presence of blasts at very low thresholds (between $10^{-3}\%$ and $10^{-5}\%$). The detection and the quantification of MRD in AML has been widely studied as a predictive and prognostic tool [18, 23, 35, 36, 16, 24]. Molecular biology and flow cytometry are the main two techniques currently available to achieve informative sensitivity levels for anticipating the disease’s progression. While molecular biology techniques for the assessment of MRD were the first to be developed and have a recognized decision-making value [32, 5], it has limited applicability (only available for 40% of the patients) due to the lack of sufficiently specific molecular targets in 60% of patients.
@@
The current preferred method for identifying cells is the so-called manual gating. It is based on the successive projection of the point cloud of the cytometry measurements according to two selected parameters. The experts then partition the space according to the two selected parameters, visually identifying regions of high density. In the diagnosis of AML, the process of manual gating is well-suited and largely employed to the detection of blasts (tumor cells) which invade the body to great amounts (at thresholds of $10^{11}$ to $10^{12}$), allowing easy identification of the tumor cell subpopulation. However, it presents a few shortcomings: it is highly dependent on the biologist’s expertise and it is extremely time-consuming. Moreover, the significant variability in flow cytometry data is a limiting factor. This variability is induced both by the biological heterogeneity between the samples and by technical details such as the cytometer performance and the data acquisition parameters. Finally, because of the very low thresholds of MRD, the identification based on manual gating of cellular subpopulation of blasts in flow cytometry data during follow-up of patients treated with chemotherapy is made tricky, and even unreasonable for biologists.
@@
For a given patient, a single dataset of flow cytometry is usually made of measurements using 10 to 20 parameters from hundreds of thousands of cells. This motivates the need to develop dimensionality reduction tools for representing variations of interest within datasets acquired from an individual patient and across a population of patients using multiple flow cytometry measurements (FCM). To overcome the shortcomings of a manual approach to data gating in FCM, the development of automatic inference methods is currently a very active field of research in computational statistics and bioinformatics. Several methods have been introduced to automate the gating of cytometry data [41]. One can cite unsupervised algorithms such as FlowMeans [1], based on K-means, Cytometree [17], based on binary trees, or [22] based on Dirichlet process mixtures. There also exist supervised algorithms such as DeepCyTOF [25] or flowlearn [26]. Regarding MRD detection for AML, [38] proposes a method using FlowSOM [37], a visualization and clustering method based on self-organising maps. In [29], the authors also use FlowSOM to detect MRD for chronic lymphocytic leukemia.
@@
Optimal Transport (OT) provides a versatile and powerful framework for addressing a diverse array of problems involving point cloud modeled as discrete probability measures. In the last decade, the use of computational OT [31] in data science and the related notion of Wasserstein distance have gained an increasing interest in statistics (see e.g. [8, 30] for recent overviews), in particular for dimension reduction of a set of probability measures using the notion of Wasserstein Principal Component Analysis (PCA) [7, 13]. OT has also proved useful for understanding various phenomena in biology [6, 12, 20, 34]. For the purpose of automatic gating from flow cytometry data, OT has been previously leveraged for quantifying the distance between two point clouds, taking into account the geometry of the data in high-dimensional spaces, giving rise two OT-based algorithms: optimalFlow [19] proposes a method for supervised gating, and CytOpt [21] estimates the different cell proportions of a population from cytometry measurements.
@@ 1.1 Contributions
We explore a statistical methodology based on computational OT for embedding into a low-dimensional linear space a set of flow cytometry measurements (FCM) modeled as high-dimensional probability distributions. Our approach is suitable for the study of datasets that consist of multiple FCM from various patients with possibly several measurements for a given patient (e.g.  acquired at different time or from different cytometers). Using the notion of Wasserstein barycenter [2], we justify the use of the K-means algorithm for dimensionality reduction of multiple large-scale point clouds through mean measure quantization by merging all the data into a single point cloud. This quantization step allows to reduce the size of the support of the observed data viewed as high-dimensional discrete probability measures. The resulting low-dimensional support quantized probability measures are then embedded into a linear space using either Wasserstein PCA [7, 13] through linearized OT [11, 40] or log-ratio PCA of compositional data [3]. This enables the analysis of intra and inter-patients FCM variability. Using a publicly available dataset and a dataset provided by Bordeaux University Hospital, this approach is first shown to achieve a meaningful clustering of patients according to their FCM. We also demonstrate the benefits of our methodology over the popular kernel mean embedding (KME) technique [28] for statistical learning from multiple high-dimensional probability distributions. Secondly, our approach is applied to the detection of MRD in AML by characterizing, in this low-dimensional representation, the patients with significant MRD based on a comparative analysis of FCM data at diagnosis and follow-up after chemotherapy treatment. In particular, our OT-based approach is found to provide a relevant and informative two-dimensional representation of the results based on the FlowSom algorithm, a state-of-the-art method for the detection of MRD in AML using FCM.
@@ 2 Materials and methods
2.1 DataML study
A first set of FCM for the statistical study conducted in this paper is provided by the Hematology Laboratory of Bordeaux University Hospital (Hôpital Haut-Lévêque). The dataset was collected from patients recruited between October 2019 and December 2022, diagnosed with acute myeloid leukemia (AML). They received intensive chemotherapy treatment and achieved complete remission after induction. This database received a favorable opinion from CCTIRS on April 2015 (file number 15.319) and from CNIL on October 2015 (authorization request no. 915285). It is also covered by the MR-004 of Bordeaux University Hospital. The Research Ethics Committee of the Bordeaux University Hospital certified that the project received a favorable opinion under the reference number CER-BDX 2024-164.
From this study, we have kept 182 patients and 22 normal bone marrow. For each patient, two or three FCM are available: one diagnosis ($\sim$50000 cells) and one or two follow-ups (up to 500000 cells). In total, we have 512 FCM. For each follow-up, we dispose of its associated MRD estimation with the FlowSOM method. A molecular biology MRD measure is also given when available (108 out of 314 follow-ups), which we consider as ground truth. There are 11 markers: 2 are structural (FS INT, SS INT) and 9 are protein expressions (CD34, CD13, CD38, CD7, CD33, CD56, CD117, DR, CD45). Hence, the datasets consist of multiple point clouds in dimension $d=11$. We represent the $N=512$ FCM as high-dimensional discrete probability measures to leverage OT statistical learning tools. For $1\leq i\leq N$, we represent the $i$-th dataset $X^{i}=(X^{i}_{1},\cdots,X^{i}_{m_{i}})\in(\mathbb{R}^{d})^{m_{i}}$ with $m_{i}$ the number of cells in this dataset, as a discrete measure $\mu^{i}=\frac{1}{m_{i}}\sum_{j=1}^{m_{i}}\delta_{X_{j}^{i}}$.
@@ 2.3 The FlowSOM method
The method in [38] to evaluate MRD in a follow-up of a patient undergoing treatment first consists in merging three cytometry entry files: a normal bone marrow (from an arbitrary, non-diseased patient), the diagnosis of the patient, and his follow-up. The authors in [38] then perform a FlowSOM on the merged files, the last step of which results in an MST. They visualize the weights on the MST for each entry file. By comparing the three MST globally, they are able to identify the Nodes of Interests (NoIs) that are decisive in assessing the MRD present in the measurements. For instance, a NoI must have a certain percentage of blasts in the diagnosis. MRD is considered as being positive in the follow-up if the percentage of cells in the NoIs in the follow-up sample is at least twice that of the same node in the normal bone marrow. It is reported in [38] that this method shows 80% concordance with molecular biology, that is with the ground truth results.
@@ 2.4.2 Dimension reduction via mean measure quantization
A first contribution of our work is to use the OT framework to justify a classical pre-processing step in flow cytometry data analysis: summarizing spatial distribution of cells for each FCM sample into a small number $K$ of common $d$-dimensional points. Indeed, the dimensionality of the data (i.e. the number of cells per FCM sample) is prohibitive for the application of any statistical analysis based on computational OT on point clouds. For example, in the FlowSOM method [38], the authors read the information of the presence of blasts in the size of the clusters generated by the SOM for each dataset.
@@ 2.6 Performance evaluation
For some cytometry datasets of the DataML study, ground truth labels are available with molecular biology, giving a percentage linked to quantities expressing the presence of blasts. MRD is then defined as being positive when this value is higher than $10^{-4}\%$ and negative otherwise. We will then refer to positive and negative follow-ups respectively. On the other hand, the FlowSOM method also provides a measure of MRD, which is related to the percentage of blasts in the bone marrow. In cytometry, the threshold for positivity is defined as $10^{-3}\%$.
@@ 3.2.1 Representation of a patient’s data
The first step of our analysis is to perform a K-means on the merged DataML dataset. This results in $N$ measures $\nu^{i}=\sum_{k=1}^{K}a^{i}_{k}\delta_{x_{k}}$ with a common support $\{x_{1},\cdots,x_{K}\}\subset\mathbb{R}^{d}$. To compare ourselves with FlowSOM [37], we propose to visualize the output of K-means with a tree. We construct a graph whose nodes are the K-means centers $x_{1},\ldots,x_{K}$. Each pair of nodes $x_{i}$ and $x_{j}$, is connected by an edge with weight corresponding to the distance $\|x_{i}-x_{j}\|^{2}$. We finally compute a tree by outputting a minimum spanning tree (MST) from this graph. One can then visualize the MST corresponding to different data of one patient in Figure 3. The size of the nodes vary from one MST to another, as it represents the number of cells in the corresponding clusters. Comparing a MST with the common MST reveals biological specificities. In the case of Figure 3, we compare the diagnosis of the patient, two of their follow-ups, and a Normal Bone Marrow (NBM). It is interesting to notice that the node 0 holds quite some mass in the diagnosis and almost none in the NBM. One could speculate that this node is a cluster of blasts. This assumption is also supported by the fact that node 0 has more mass in the second follow-up where molecular MRD is higher, than in the first follow-up where MRD is lower.
Figure 3: Minimum Spanning Tree for the diagnosis and two follow-up measurements of one patient of the DataML dataset, and one normal bone marrow.
@@ 3.2.2 PCA
In Figure 5, we visualize the amount of MRD in the follow-ups for the K-Means + LinW2 PCA method, where the size of the follow-up points represents the quantity of molecular biology for the patient at the time that the sample was collected. We observe that high MRD follow-ups are pulled towards the set of diagnosis. We also visualize the MRD prediction of the FlowSOM method in Figure 6, where we observe the same effect: high predicted MRD points lay in the diagnosis set whereas lower predicted MRD is closer to normal bone marrow points. Our low-dimensional representation of multi-patient FCM datasets is therefore consistent with the results of the FlowSOM method on the level of estimated MRD by the approach proposed in [38].
Figure 4: Two-dimensional representations of the DataML datasets using PCA on different embeddings : (top) K-Means + LinW2, (middle) K-Means + comp, (bottom) KME. Each dot represents a FCM sample. The colors encore the nature of the data : (red) diagnostic, (purple) normal bone marrow, (yellow) positive follow-up, (green) negative follow-up, (grey) no information on the MRD.
Figure 5: Two-dimensional representation of the DataML datasets using PCA on the K-Means + LinW2 embedding. The size of the follow-ups is scaled with a logicle transformation of MRD measured with molecular biology ($\log(y)$ when $y\geq 1$, $y$ else).
Figure 6: Two-dimensional representation of the DataML datasets using PCA on the K-Means + LinW2 embedding. The size of the follow-ups is scaled to the predicted MRD by the FlowSOM method.

### v2

# Low dimensional representation of multi-patient flow cytometry datasets using optimal transport for measurable residual disease detection in leukemia

@@ Abstract
Representing and quantifying Measurable Residual Disease (MRD) in Acute Myeloid Leukemia (AML), a type of cancer that affects the blood and bone marrow, is essential in the prognosis and follow-up of AML patients. As traditional cytological analysis cannot detect leukemia cells below 5%, the analysis of flow cytometry dataset is expected to provide more reliable results. In this paper, we explore statistical learning methods based on optimal transport (OT) to achieve a relevant low-dimensional representation of multi-patient flow cytometry measurements (FCM) datasets considered as high-dimensional probability distributions. Using the framework of OT, we justify the use of the K-means algorithm for dimensionality reduction of multiple large-scale point clouds through mean measure quantization by merging all the data into a single point cloud. After this quantization step, the visualization of the intra and inter-patients FCM variability is carried out by embedding low-dimensional quantized probability measures into a linear space using either Wasserstein Principal Component Analysis (PCA) through linearized OT or log-ratio PCA of compositional data. Using a publicly available FCM dataset and a FCM dataset from Bordeaux University Hospital, we demonstrate the benefits of our approach over the popular kernel mean embedding technique for statistical learning from multiple high-dimensional probability distributions. We also highlight the usefulness of our methodology for low-dimensional projection and clustering patient measurements according to their level of MRD in AML from FCM. In particular, our OT-based approach allows a relevant and informative two-dimensional representation of the results of the FlowSom algorithm, a state-of-the-art method for the detection of MRD in AML using multi-patient FCM.
@@ 1 Introduction
Acute Myeloid Leukemia (AML) is a cancer of the blood and bone marrow when the development of blood stem cells (immature cells) into mature cells is altered, and the number of these immature cells, called myeloid blasts, increases. The diagnosis of AML consists in the detection of blasts which invade the bone marrow to great amounts, and it is typically assessed with the cytological analysis of the patient’s bone marrow. The goal of the chemotherapy treatment is therefore to eradicate the blasts population: 80% of patients reach complete remission after one cycle of chemotherapy, however 50% relapse [40]. The definition of complete remission states that the bone marrow should contain less than 5% of blasts. However, detecting smaller numbers of persisting residual leukemic cells, called Measurable Residual Disease (MRD), is critical for disease monitoring and relapse prognosis. This motivates the need to catch the presence of blasts at very low thresholds (between 0.1% and 0.001%). The detection and the quantification of MRD in AML has been widely studied as a predictive and prognostic tool [18, 23, 37, 38, 16, 24]. Molecular biology and flow cytometry are the main two techniques currently available to achieve informative sensitivity levels for anticipating the disease’s progression. While molecular biology techniques for the assessment of MRD were the first to be developed and have a recognized decision-making value [34, 5], it has limited applicability (only available for 40% of the patients) due to the lack of sufficiently specific molecular targets in 60% of patients.
@@
The current preferred method for identifying cells is the so-called manual gating. It is based on the successive projection of the point cloud of the cytometry measurements according to two selected parameters. The experts then partition the space according to the two selected parameters, visually identifying regions of high density. In the diagnosis of AML, the process of manual gating is well-suited and largely employed to the detection of blasts (tumor cells) which invade the body to great amounts (at thresholds of $10^{11}$ to $10^{12}$), allowing easy identification of the tumor cell subpopulation. However, it presents a few shortcomings: it is highly dependent on the biologist’s expertise and it is extremely time-consuming. Moreover, the significant variability in flow cytometry data is a limiting factor. This variability is induced both by the biological heterogeneity between the samples and by technical details such as the cytometer performance and the data acquisition parameters. Finally, because of the very low thresholds of MRD, the identification based on manual gating of cellular subpopulation of blasts in flow cytometry data during follow-up of patients treated with chemotherapy is made tricky.
@@
For a given patient, a single dataset of flow cytometry is usually made of measurements using 10 to 20 parameters from hundreds of thousands of cells. This motivates the need to develop dimensionality reduction tools for representing variations of interest within datasets acquired from an individual patient and across a population of patients using multiple flow cytometry measurements (FCM). To overcome the shortcomings of a manual approach to data gating in FCM, the development of automatic inference methods is currently a very active field of research in computational statistics and bioinformatics. Several methods have been introduced to automate the gating of cytometry data [43]. One can cite unsupervised algorithms such as FlowMeans [1], based on K-means, Cytometree [17], based on binary trees, or [22] based on Dirichlet process mixtures. There also exist supervised algorithms such as DeepCyTOF [25] or flowlearn [26]. Regarding MRD detection for AML, [40] proposes a method using FlowSOM [39], a visualization and clustering method based on self-organising maps. In this paper, we will denote MRD-FlowS the MRD measure calculated by this method. On the other hand, MRD-BioM will designate the MRD value assessed with molecular biology. In [31], the authors also use FlowSOM to detect MRD for chronic lymphocytic leukemia.
@@
Optimal Transport (OT) provides a versatile and powerful framework for addressing a diverse array of problems involving point cloud modeled as discrete probability measures. In the last decade, the use of computational OT [33] in data science and the related notion of Wasserstein distance have gained an increasing interest in statistics (see e.g. [7, 32] for recent overviews), in particular for dimension reduction of a set of probability measures using the notion of Wasserstein Principal Component Analysis (PCA) [8, 13]. OT has also proved useful for understanding various phenomena in biology [6, 12, 20, 36]. For the purpose of automatic gating from flow cytometry data, OT has been previously leveraged for quantifying the distance between two point clouds, taking into account the geometry of the data in high-dimensional spaces, giving rise to two OT-based algorithms: optimalFlow [19] proposes a method for supervised gating, and CytOpt [21] estimates the different cell proportions of a population from cytometry measurements. More recently, [27] introduced a method based on linear optimal transport for the classification of point clouds and showed its applications to flow cytometry datasets from AML patients.
@@ 1.1 Contributions
We explore a statistical methodology based on computational OT for embedding into a low-dimensional linear space a set of flow cytometry measurements (FCM) modeled as high-dimensional probability distributions. Our approach is suitable for the study of datasets that consist of multiple FCM from various patients with possibly several measurements for a given patient (e.g.  acquired at different time or from different cytometers). Using the notion of Wasserstein barycenter [2], we justify the use of the K-means algorithm for dimensionality reduction of multiple large-scale point clouds through mean measure quantization by merging all the data into a single point cloud. This quantization step allows to reduce the size of the support of the observed data viewed as high-dimensional discrete probability measures. The resulting low-dimensional support quantized probability measures are then embedded into a linear space using either Wasserstein PCA [8, 13] through linearized OT [11, 42, 29, 27] or log-ratio PCA of compositional data [3]. This enables the analysis of intra and inter-patients FCM variability. Using a publicly available dataset and a dataset provided by Bordeaux University Hospital, this approach is first shown to achieve a meaningful clustering of patients according to their FCM. We also demonstrate the benefits of our methodology over the popular kernel mean embedding (KME) technique [30] for statistical learning from multiple high-dimensional probability distributions. Secondly, we study applications to the assessment of MRD in AML. In particular, our OT-based approach is found to provide a relevant and informative two-dimensional representation of the results obtained with the FlowSom algorithm (MRD-FlowS), a robust method with growing adoption in the field for the detection of MRD in AML using FCM. This representation could be a valuable tool to help biologists confirm or infirm the relapse prognosis of AML patients.
@@ 2 Materials and methods
2.1 DATAML-Bordeaux study
A first set of FCM for the statistical study conducted in this paper is provided by the Hematology Laboratory of Bordeaux University Hospital (Hôpital Haut-Lévêque). The dataset was collected from patients recruited between October 2019 and December 2022, diagnosed with acute myeloid leukemia (AML). They received intensive chemotherapy treatment and achieved complete remission after induction. This database is covered by the MR-004 of Bordeaux University Hospital. The Research Ethics Committee of the Bordeaux University Hospital certified this project and received a favorable opinion under the reference number CER-BDX 2024-164.
From this study, we have kept 182 patients and 22 normal bone marrow (NBM). For each patient, two or three FCM are available: one diagnosis ($\sim$50000 cells) and one or two follow-ups (up to 500000 cells). In total, we have 512 FCM. For each follow-up, we dispose of its associated MRD-FlowS estimation with the FlowSOM method. The MRD-BioM measure is also given when available (108 out of 314 follow-ups), which we consider as ground truth. There are 11 markers: 2 are structural (FS INT, SS INT) and 9 are protein expressions (CD34, CD13, CD38, CD7, CD33, CD56, CD117, HLA-DR, CD45). Hence, the datasets consist of multiple point clouds in dimension $d=11$. We represent the $N=512$ FCM as high-dimensional discrete probability measures to leverage OT statistical learning tools. For $1\leq i\leq N$, we represent the $i$-th dataset $X^{i}=(X^{i}_{1},\cdots,X^{i}_{m_{i}})\in(\mathbb{R}^{d})^{m_{i}}$ with $m_{i}$ the number of cells in this dataset, as a discrete measure $\mu^{i}=\frac{1}{m_{i}}\sum_{j=1}^{m_{i}}\delta_{X_{j}^{i}}$.
@@ 2.3 The FlowSOM method
The method in [40] to evaluate MRD in a follow-up of a patient undergoing treatment first consists in merging three cytometry entry files: a normal bone marrow (from an arbitrary, healthy donor), the diagnosis of the patient with AML, and his follow-up. The authors in [40] then perform a FlowSOM on the merged files, the last step of which results in an MST. They visualize the weights on the MST for each entry file. By comparing the three MST globally, they are able to identify the Nodes of Interests (NoIs) that are decisive in assessing the MRD present in the measurements. It is reported in [40] that this method shows 80% concordance with MRD-Bio, that is with the ground truth results. More precisely, MRD-FlowS yields 85% specificity and 69% sensitivity compared to MRD-BioM.
@@ 2.4.2 Dimension reduction via mean measure quantization
A first contribution of our work is to use the OT framework to justify a classical pre-processing step in flow cytometry data analysis: summarizing spatial distribution of cells for each FCM sample into a small number $K$ of common $d$-dimensional points. Indeed, the dimensionality of the data (i.e. the number of cells per FCM sample) is prohibitive for the application of any statistical analysis based on computational OT on point clouds. For example, in the FlowSOM method [40], the authors read the information of the presence of blasts in the size of the clusters generated by the SOM for each dataset, focusing on clusters corresponding to cells compatible with blasts in a CD45/SS graph.
@@ 2.6 Performance evaluation
For some cytometry datasets of the DATAML-Bordeaux study, ground truth labels are available with the measures (MRD-BioM), giving a percentage linked to quantities expressing the presence of blasts. MRD-BioM is then defined as being positive when its value is higher than $0.1\%$ and negative otherwise. We will then refer to positive and negative follow-ups respectively. On the other hand, the FlowSOM method also provides a measure of MRD, MRD-FlowS, which is related to the percentage of blasts in the bone marrow. In cytometry, the threshold for positivity is also defined as $0.1\%$.
@@ 3.2.1 Representation of a patient’s data
The first step of our analysis is to perform a K-means on the merged DATAML-Bordeaux dataset. This results in $N$ measures $\nu^{i}=\sum_{k=1}^{K}a^{i}_{k}\delta_{x_{k}}$ with a common support $\{x_{1},\cdots,x_{K}\}\subset\mathbb{R}^{d}$. To compare ourselves with FlowSOM [39], we propose to visualize the output of K-means with a tree. We construct a graph whose nodes are the K-means centers $x_{1},\ldots,x_{K}$. Each pair of nodes $x_{i}$ and $x_{j}$ is connected by an edge with its weight corresponding to the distance $\|x_{i}-x_{j}\|^{2}$. We then compute a minimum spanning tree (MST) from this graph. One can visualize this MST corresponding to different data of one patient in Figure 3. The size of the nodes vary from one MST to another, as it represents the number of cells in the corresponding clusters. Comparing a MST with the common MST reveals biological specificities. In the case of Figure 3, we compare the diagnosis of the patient, two of their follow-ups, and a NBM. It is interesting to notice that the node at the center of the tree holds quite some mass in the diagnosis and almost none in the NBM. One could speculate that this node is a cluster of blasts. This assumption is also supported by the fact that this node has more mass in the second follow-up where the MRD-BioM value is higher, than in the first follow-up where the MRD-BioM value is lower.
Figure 3: Minimum Spanning Tree for the diagnosis and two follow-up measurements of one patient of the DATAML-Bordeaux dataset, and one normal bone marrow.
@@ 3.2.2 PCA
In Figure 5, we visualize the amount of MRD in the follow-ups for the K-Means + LinW2 PCA method, where the size of the follow-up points represents the MRD-BioM value for the patient at the time that the sample was collected. We observe that high MRD-BioM follow-ups are pulled towards the set of diagnosis. Regarding the negative follow-ups, these are evidently more directed towards NBM and clearly separated from the set of diagnosis. We also visualize the MRD-FlowS predictions with the FlowSOM method in Figure 6, where we observe the same effect: high MRD-FlowS points lay in the diagnosis set whereas lower MRD-FlowS are closer to normal bone marrow points. Our low-dimensional representation of multi-patient FCM datasets is therefore consistent with the values of MRD-FlowS.
Figure 4: Two-dimensional representations of the DATAML-Bordeaux datasets using PCA on different embeddings : (top) K-Means + LinW2, (middle) K-Means + comp, (bottom) KME. Each dot represents a FCM sample. The colors encore the nature of the data : (red) diagnostic, (purple) normal bone marrow, (yellow) positive follow-up, (green) negative follow-up, (grey) no information on the MRD-BioM.
Figure 5: Two-dimensional representation of the DATAML-Bordeaux datasets using PCA on the K-Means + LinW2 embedding. The size of the follow-ups is scaled with a logicle transformation of the MRD-BioM values ($\log(y)$ when $y\geq 1$, $y$ else).
Figure 6: Two-dimensional representation of the DATAML-Bordeaux datasets using PCA on the K-Means + LinW2 embedding. The size of the follow-ups is scaled to the values of MRD-FlowS.